how many grams of NH4NO3 are in 125mL of 1.40M solution?
Answer #1

molarity = (W /MW) (1000 / V in mL)

molarity = 1.40 M

W = ?

MW = 80.04 g/mol

V = 125 mL

1.40 = (W / 80.04) (1000 / 125)

W / 80.04 = 0.175

W = 14.0 g

mass of NH4NO3 required = 14.0 g